# 2cm2cm1cm1cm

### v4

@@ 3.2 Continuous Flow of Bogoliubov Transformations
We need to prove that the corresponding non-autonomous evolution equation is well-established with time-dependent generators (57). Given $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$, Proposition 3.4 proves that quadratic operators like the infinitesimal generators $\mathrm{G}_{t}$, $t\in\mathbb{R}_{0}^{+}$, all defined on the same (dense) domain
@@
[…]
@@
for all $s,t\in\mathbb{R}_{0}^{+}$. Note that these derivatives [^19] The derivatives $\partial_{t}$ and $\partial_{s}$ on the borderline $t=0$ or $s=0$ have to be understood as right derivatives. follow from $D_{t}=-D_{t}^{\top}$ and the equalities
@@
$U_{t,s}=\left(\begin{array}[]{cc}\mathbf{u}_{t,s}&\mathbf{v}_{t,s}\\ \overline{\mathbf{v}}_{t,s}&\overline{\mathbf{u}}_{t,s}\end{array}\right)\ ,\qquad s,t\in\mathbb{R}_{0}^{+}\ ,$
@@
There is a small inconsistency or typo in the first assertion of [6, Lemma 71] as one can see from its proof (which is correct). [6, Lemma 71] done in the bosonic case has be analogous to Proposition 3.6.
@@ 3.3 Solution to the Brockett-Wegner Flow on Quadratic Hamiltonians
The continuous flow of Bogoliubov transformations used here is based on an elliptic flow (191) on operators acting on the one-particle Hilbert space. This flow is not studied here, but in [24], some important results of which are shortly explained in Section A. Therefore, the kernel of our proofs on the diagonalization of quadratic Hamiltonians lies on the one-particle Hilbert space.
@@
In fact, our aim here is not only to obtain a diagonalization of quadratic Hamiltonians, but also to highlight the Brockett-Wegner flow [19, 20, 21, 22] by given another rigorous solution (and even its asymptotics) of this (usually) formal method for an unbounded case. In particular, trying to diagonalize quadratic Hamiltonians by adapting the method developed in [8] for the bosonic case may be efficient, but it would definitively suppress our second aim here, namely the study of the Brockett-Wegner flow and ODEs on operator spaces like the elliptic flow (191). Indeed, differential equations on non-commutative algebras of operators (like the Brockett-Wegner flow [19, 20, 21, 22]) are very little developed in Mathematics (in particular for infinite-dimensional spaces). Hence, any mathematical result on relevant differential equations for operators is useful and possibly liminal. This is one important objective of [24] and the present paper, the diagonalization of quadratic Hamiltonians under much more general conditions being one salient application achieved via asymptotics results of Section 3.5. Moreover, note that comparatively to [8] (done in the bosonic case), we obtain here some expressions for the diagonal form as integrals of elements of the elliptic flow, and in some cases a very explicit expressions (see (16) and (17)).
@@
for some $\0\lambda\in\mathbb{R}\backslash\{0\}$. The proof is the same as the one done in various lemmata of [6, Section VI.2], except one has to replace [6, Lemma 65] with Corollary 3.2. For the reader’s convenience, we give below the important steps, but without much details.
@@ 3.4 Asymptotics of the Solution to the Brockett-Wegner Flow
on the domain $\mathcal{D}_{0}$ (5), keeping always in mind that $D$ is solution to the elliptic operator flow (191). See (35) extended to $t=\infty$. Note $\mathrm{H}_{\infty}\equiv\mathrm{H}_{\infty}^{\ast\ast}=\mathrm{H}_{\infty}^{\ast}$ is well-defined, for instance via Proposition 3.4. Up to a constant, it is the well-known second-quantization *$d\Gamma(\Upsilon_{\infty})$ (50) of $\Upsilon_{\infty}$ and its spectral properties can be directly obtained from the ones of the one-particle self-adjoint operator $\Upsilon_{\infty}$. In fact, $\mathrm{H}_{\infty}$ is $\mathrm{N}$–diagonal since it commutes with the particle number operator $\mathrm{N}$. Last but not least, it is the limit $t\rightarrow\infty$, in the strong resolvent sense, of the family $\left(\mathrm{H}_{t}\right)_{t\geq 0}$ of unitary equivalent quadratic Hamiltonians.
@@
and for any non-zero real $\0\lambda\in\mathbb{R}\backslash\{0\}$ and vector $\varphi\in\mathcal{F}$,
@@
Proof. Because of the square-integrability of the mapping $t\mapsto\|D_{t}\|_{2}$ on $\mathbb{R}_{0}^{+}$, we can extend (84) to all $t,s\in\mathbb{R}_{0}^{+}\cup\{\infty\}$ with $D_{\infty}\doteq 0$ and obviously deduce (91). To prove next that $\left(\mathrm{H}_{t}\right)_{t\geq 0}$ converges to $\mathrm{H}_{\infty}$ in the strong resolvent sense, observe first that, for any non-zero $\0\lambda\in\mathbb{R}\backslash\{0\}$ and $C\in\mathbb{R}^{+}$, there is $s_{\lambda}\in\mathbb{R}_{0}^{+}$ such that
@@
because, by assumption, the Hilbert-Schmidt norm $\|D_{t}\|_{2}$ vanishes when $t\rightarrow\infty$. If the mapping $t\mapsto\|D_{t}\|_{2}$ is square-integrable on $\mathbb{R}_{0}^{+}$ then Equations (82)–(83) are satisfied for $t=\infty$ and sufficiently large $s\in\mathbb{R}_{0}^{+}$. As a consequence, using the resolvent identity, for any $\0\lambda\in\mathbb{R}\backslash\{0\}$, sufficiently large initial time $s\in\mathbb{R}_{0}^{+}$ and $t\in\left[s,\infty\right)$,
@@
for any $\0\lambda\in\mathbb{R}\backslash\{0\}$, thanks to (82)–(83) for $t=\infty$ and (91). The domain $\mathcal{D}(\mathrm{N})$ is dense in $\mathcal{F}$ and both resolvents $(\mathrm{H}_{\infty}+i\lambda\mathbf{1})^{-1}$ and $(\mathrm{H}_{t}+i\lambda\mathbf{1})^{-1}$ in (94) are bounded operators for any $t\in\mathbb{R}_{0}^{+}$ and for any fixed $\0\lambda\in\mathbb{R}\backslash\{0\}$. As a consequence, the strong resolvent convergence follows from (94) and a standard continuity argument.
@@ 3.5 Diagonalization of Quadratic Hamiltonians
Following Brockett-Wegner’s strategy – or Theorem 3.8 and Proposition 3.9 in our example – the diagonalization of quadratic Hamiltonians $\mathrm{H}_{0}$ (4) via a Bogoliubov $\mathbf{u}$$\mathbf{v}$ transformation should be achieved by performing the limits $t\rightarrow\infty$ of bounded operator families $\left(\mathrm{U}_{t,s}\right)_{t,s\in\mathbb{R}_{0}^{+}}$ (Theorems 3.5) and $\left(\mathbf{u}_{t,s}\right)_{t,s\in\mathbb{R}^{+}}$, $\left(\mathbf{v}_{t,s}\right)_{t,s\in\mathbb{R}_{0}^{+}}$ (see (66)–(67) and Proposition 3.6). Recall indeed that the limit $t\rightarrow\infty$ of the solution $\left(\mathrm{H}_{t}\right)_{t\in\mathbb{R}_{0}^{+}}$ to the Brockett-Wegner flow (30) leads to the $\mathrm{N}$–diagonal quadratic Hamiltonian $\mathrm{H}_{\infty}$ (90), thanks to Proposition 3.9. This refers to a quasi $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians. To get a true $\mathrm{N}$–diagonalization we prove that $\mathrm{H}_{\infty}$ is unitary equivalent to any $\mathrm{H}_{t}$, $t\in\mathbb{R}_{0}^{+}$.
@@
$\mathrm{U}_{\infty,\infty}\varphi\doteq\underset{t\rightarrow\infty}{\lim}\underset{s\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}=\underset{s\rightarrow\infty}{\lim}\underset{t\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}=\varphi\ .$ (101)
In particular, $\mathrm{U}_{\infty,\infty}=\mathbf{1}$. It is also straightforward to verify from (97) and (99) that $\left(\mathrm{U}_{\infty,s}\right)_{s\in\mathbb{R}_{0}^{+}}$ and $\left(\mathrm{U}_{t,\infty}\right)_{t\in\mathbb{R}_{0}^{+}}$ are strongly continuous in $s$ and $t$, respectively, while $\mathrm{U}_{t,s}=\mathrm{U}_{t,x}\mathrm{U}_{x,s}$ for any $x,t,s\in\mathbb{R}_{0}^{+}\cup\left\{\infty\right\}$. In the same way, one easily checks that they are family of unitary operators with $\mathrm{U}_{\infty,s}^{\ast}=\mathrm{U}_{s,\infty}$ for any $s\in\mathbb{R}_{0}^{+}$. Moreover, using the triangle inequality, one gets that, for any $\0\epsilon\in\mathbb{R}\backslash\{0\}$ with sufficiently small $\left|\epsilon\right|>0$, all times $s,t\in\mathbb{R}^{+}$ and any vector $\varphi\in\mathcal{D}\left(\mathrm{N}^{2}\right)\subseteq\mathcal{D}\left(\mathrm{N}\right)\subseteq\mathcal{D}\left(\mathrm{G}_{s}\right)$,
@@
Note that Theorem 3.10 for the bosonic case is only partially proven in [6], even though it can also be shown in this situation. Remark additionally that $n\in\mathbb{N}$ iterations of the formal computation (61) to compute the $n$-folds muticommutators $[\mathrm{N},[\mathrm{N},\cdots,[\mathrm{N},\mathrm{G}_{t}]\cdots]$ together with Corollary 3.2 yield
@@
3.6 Shale-Stinespring-like Condition for Generators of Bogoliubov Transformations
As explained in Section 2.6, Bach, Lieb and Solovej [15] introduced a general definition of quadratic Hamiltonians on fermionic Fock spaces. This refers to Definition 2.10. It leads to a strongly continuous one-parameter group of Bogoliubov transformations, as explained in the proof of Lemma 2.12. In this section we study such kind of groups and discuss their generators in relation with the Shale-Stinespring condition (see Theorem 2.8).
We take $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and recall that a Bogoliubov transformation is a unitary operator $U\in\mathcal{B}(\mathcal{H})$ such that $U\mathfrak{A=A}U$, as explained in Section 2.4. We thus start by relating these properties with the generator of a group of Bogoliubov transformations:
@@ 3.6 Shale-Stinespring-like Condition for Generators of Bogoliubov Transformations
Proof. Any bounded operator $U$ acting on $\mathcal{H}$ can always be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ by
$U=\left(\begin{array}[]{cc}P_{\mathfrak{h}}UP_{\mathfrak{h}}&P_{\mathfrak{h}}UP_{\mathfrak{h}}^{\perp}\\ P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}&P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}^{\perp}\end{array}\right)$ (108)
with $P_{\mathfrak{h}}$ being the orthogonal projection on $\mathfrak{h}\oplus\{0\}$ and $P_{\mathfrak{h}}^{\perp}\doteq\mathbf{1}-P_{\mathfrak{h}}$. (By a slight abuse of notation, we can use in an appropriate way the identification $\mathfrak{h}\oplus\{0\}\equiv\mathfrak{h}$ and $\{0\}\oplus\mathfrak{h}\equiv\mathfrak{h}$ to see $P_{\mathfrak{h}}UP_{\mathfrak{h}}$, $P_{\mathfrak{h}}UP_{\mathfrak{h}}^{\perp}$, $P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}$ and $P_{\mathfrak{h}}^{\perp}UP_{\mathfrak{h}}^{\perp}$ as four operators acting on $\mathfrak{h}$.) If $U$ is a Bogoliubov transformation, i.e., $UU^{\ast}=U^{\ast}U=\mathbf{1}$ and $U\mathfrak{A=A}U$ then, using (21) together with $\overline{X}\doteq\mathcal{C}X\mathcal{C}$ (Remark 2.1) one computes that the above representation can be reduced to
$U=\left(\begin{array}[]{cc}\mathbf{u}&\mathbf{v}\\ \overline{\mathbf{v}}&\overline{\mathbf{u}}\end{array}\right)\ ,\qquad\mathbf{u},\mathbf{v}\in\mathcal{B}\left(\mathfrak{h}\right)\ .$ (109)
Applied to the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations, we get Assertion (i). It remains to represent the self-adjoint Hamiltonian $h$ leading to the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$. A priori, one can represent $h$ on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ like (108) for $U=h$, but as explained after Equation (38), there are nontrivial domain issues. To have some control on the domain of the corresponding operators, we first use Lemma 3.13 to deduce that $h$ is a self-dual Hamiltonian, i.e., $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$. Secondly, recall that basis projections are orthogonal projections $P\in\mathcal{B}(\mathcal{H})$ satisfying $\mathfrak{A}P\mathfrak{A}=P^{\bot}\doteq\mathbf{1}_{\mathcal{H}}-P$. If $\dim\ker h\in 2\mathbb{N}\cup\{\infty\}$ and $h=-\mathfrak{A}h\mathfrak{A}$, then [18, Lemma 4.8 and its proof] leads to the existence of a pair of basis projections $P$ and $P^{\bot}\doteq\mathbf{1}-P$ with range $\mathfrak{h}_{P}$ and $\mathfrak{h}_{P^{\bot}}$, respectively, such that $h$ can be represented on $\mathcal{H}=\mathfrak{h}_{P}\oplus\mathfrak{h}_{P^{\bot}}$ by
@@
where $\mathbb{A}=\mathbb{A}^{\ast}$ is a (possibility unbounded) self-adjoint operator with domain $\mathcal{D}\left(\mathbb{A}\right)\subseteq\mathfrak{h}_{P}$. Even in the finite dimension situation, one can use the identification $\mathfrak{h}_{P}\equiv\mathfrak{A}\mathfrak{h}_{P}=\mathfrak{h}_{P^{\bot}}$ and see $-\mathbb{A}^{\top}$ as an operator acting on $\mathfrak{h}_{P^{\bot}}$. We meanwhile observe from [18, Lemma 4.8] that
@@
In the above matrix representation of $h$, note that we used $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$ (Remark 2.1), which yields $\mathbb{A}^{\top}=\overline{\mathbb{A}}=\mathcal{C\mathbb{A}C}$. By [18, Lemma 3.6], there is a Bogoliubov transformation $U$ such that $h$ can be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ via the following product
$h=U^{\ast}\left(\begin{array}[]{cc}\mathbb{A}&0\\ 0&-\mathbb{A}^{\top}\end{array}\right)U\ .$
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. Let $c\in\mathbb{R}^{+}\cup\{\infty\}$ and assume that
@@
Observe that Assumption 1 and Condition (127) imply that $\mathcal{Y}$ must be a core for the self-adjoint operator $\Upsilon_{0}$, by [44, Proposition 6.3]. In other words, $\Upsilon_{0}$ restricted to the subspace $\mathcal{Y}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)$ is essentially self-adjoint. Note also that Lemma 3.15 can always be applied whenever the off-diagonal operator-valued coefficient $\overline{D}_{0}$ is relatively bounded with respect of the diagonal one $\Upsilon_{0}$ (see (47)). This is a direct consequence of the following lemma:
@@
From now, we restrict our focus from (possibly) fully unbounded cases to bounded off-diagonal operator-valued coefficients, i.e., $D_{0}\in\mathcal{B}(\mathfrak{h})$. It is worth mentioning that the diagonal element $\Upsilon_{0}$ is still taken as a possibly unbounded operator. In this case, using very standard arguments with Lemmata 3.15 and 3.16, we easily write all operator-valued coefficients of groups of Bogoliubov transformations via Dyson series involving operator-valued coefficients of their generator.
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations as

### v6

@@ 3.2 Continuous Flow of Bogoliubov Transformations
We need to prove that the corresponding non-autonomous evolution equation is well-posed with time-dependent generators (53). Given $D\in C(\mathbb{R}_{0}^{+};\mathcal{L}^{2}(\mathfrak{h}))$ with $D_{t}=-D_{t}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}_{0}^{+}$, Proposition 3.4 proves that quadratic operators like the infinitesimal generators $\mathrm{G}_{t}$, $t\in\mathbb{R}_{0}^{+}$, all defined on the same (dense) domain
@@
$\mathbf{u}_{t,s}^{\ast}\mathbf{u}_{t,s}+\mathbf{v}_{t,s}^{\top}\overline{\mathbf{v}}_{t,s}$ $=$ $\mathbf{1}\ ,\text{\hskip 20.44434pt}\mathbf{u}_{t,s}^{\ast}\mathbf{v}_{t,s}+\mathbf{v}_{t,s}^{\top}\overline{\mathbf{u}}_{t,s}=0\ ,$ (65)
@@
for all $s,t\in\mathbb{R}_{0}^{+}$. Note that these derivatives [^23] The derivatives $\partial_{t}$ and $\partial_{s}$ on the borderline $t=0$ or $s=0$ have to be understood as right derivatives. follow from $D_{t}=-D_{t}^{\top}$ and the equalities (62) and (63) rewritten as
@@
$\mathcal{U}_{t,s}=\left(\begin{array}[]{cc}\mathbf{u}_{t,s}&\mathbf{v}_{t,s}\\ \overline{\mathbf{v}}_{t,s}&\overline{\mathbf{u}}_{t,s}\end{array}\right)\ ,\qquad s,t\in\mathbb{R}_{0}^{+}\ ,$
@@
There is a small inconsistency or typo in the first assertion of [6, Lemma 71] as one can see from its proof (which is correct). [6, Lemma 71] done in the bosonic case has to be analogous to Proposition 3.6.
@@ 3.3 Solution to the Brockett-Wegner Flow on Quadratic Hamiltonians
The continuous flow of Bogoliubov transformations used here is based on an elliptic flow (180) on operators acting on the one-particle Hilbert space. This flow is not studied here, but in [24], some important results of which are shortly explained in Section A.1. Therefore, the kernel of our proofs on the diagonalization of quadratic Hamiltonians lies on the one-particle Hilbert space.
@@
In fact, our aim here is not only to obtain a diagonalization of quadratic Hamiltonians, but also to highlight the Brockett-Wegner flow [19, 20, 21, 22] by giving another rigorous solution (and even its asymptotics) of this (usually) formal method for an unbounded case. In particular, trying to diagonalize quadratic Hamiltonians by adapting the method developed in [8] for the bosonic case is a priori a natural idea, but it would definitively suppress our second aim here, namely the study of the Brockett-Wegner flow and ODEs on operator spaces like the elliptic flow (180). Indeed, differential equations on non-commutative algebras of operators (like the Brockett-Wegner flow [19, 20, 21, 22]) are very little developed in Mathematics (in particular for infinite-dimensional spaces). Hence, any mathematical result on relevant differential equations for operators is useful and possibly liminal. This is one important objective of [24] and the present paper, the diagonalization of quadratic Hamiltonians under much more general conditions being one salient application achieved via asymptotics results of Section 3.5. Moreover, note that comparatively to [8] (done in the bosonic case), we obtain here some expressions for the diagonal form as integrals of elements of the elliptic flow, and in some cases a very explicit expressions (see (16) and (17)).
@@
[…]
@@ 3.4 Asymptotics of the Solution to the Brockett-Wegner Flow
on the domain $\mathcal{D}_{0}$ (5), keeping always in mind that $D$ is solution to the elliptic operator flow (180). See (35) extended to $t=\infty$. Note $\mathrm{H}_{\infty}\equiv\mathrm{H}_{\infty}^{\ast\ast}=\mathrm{H}_{\infty}^{\ast}$ is well-defined, for instance via Proposition 3.4. Up to a constant, it is the well-known second-quantization $d\Gamma(\Upsilon_{\infty})$ (47) of $\Upsilon_{\infty}$ and its spectral properties can be directly obtained from the ones of the one-particle self-adjoint operator $\Upsilon_{\infty}$. In fact, $\mathrm{H}_{\infty}$ is $\mathrm{N}$–diagonal since it commutes with the particle number operator $\mathrm{N}$. Last but not least, it is the limit $t\rightarrow\infty$, in the strong resolvent sense, of the family $\left(\mathrm{H}_{t}\right)_{t\geq 0}$ of unitary equivalent quadratic Hamiltonians.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.5 Diagonalization of Quadratic Hamiltonians
Following Brockett-Wegner’s strategy – or Theorem 3.8 and Proposition 3.9 in our example – the diagonalization of quadratic Hamiltonians $\mathrm{H}_{0}$ (4) via a Bogoliubov $\mathbf{u}$$\mathbf{v}$ transformation should be achieved by performing the limits $t\rightarrow\infty$ of bounded operator families $\left(\mathrm{U}_{t,s}\right)_{t,s\in\mathbb{R}_{0}^{+}}$ (Theorems 3.5) and $\left(\mathbf{u}_{t,s}\right)_{t,s\in\mathbb{R}^{+}}$, $\left(\mathbf{v}_{t,s}\right)_{t,s\in\mathbb{R}_{0}^{+}}$ (see (62)–(63) and Proposition 3.6). Recall indeed that the limit $t\rightarrow\infty$ of the solution $\left(\mathrm{H}_{t}\right)_{t\in\mathbb{R}_{0}^{+}}$ to the Brockett-Wegner flow (30) leads to the $\mathrm{N}$–diagonal quadratic Hamiltonian $\mathrm{H}_{\infty}$ (86), thanks to Proposition 3.9. This refers to a quasi $\mathrm{N}$–diagonalization of fermionic quadratic Hamiltonians. To get a true $\mathrm{N}$–diagonalization we prove that $\mathrm{H}_{\infty}$ is unitarily equivalent to any $\mathrm{H}_{t}$, $t\in\mathbb{R}_{0}^{+}$.
@@
$\mathrm{U}_{\infty,\infty}\varphi\doteq\underset{t\rightarrow\infty}{\lim}\underset{s\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}=\underset{s\rightarrow\infty}{\lim}\underset{t\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,s}\varphi\right\}=\underset{t\rightarrow\infty}{\lim}\left\{\mathrm{U}_{t,t}\varphi\right\}=\varphi\ .$ (97)
[…]
@@
Note that Theorem 3.10 for the bosonic case is only partially proven in [6], even though it can also be shown in this situation. Remark additionally that $n\in\mathbb{N}$ iterations of the formal computation (57) to compute the $n$-folds muticommutators $\mathrm{N},[\mathrm{N},\cdots,[\mathrm{N},\mathrm{G}_{t}]\cdots]$ together with Corollary 3.2 yield
@@
4 Technical Results: BLS’s Approach to Quadratic Operators
In the previous section we prove of Theorem 2.4. In this one, we explicit the relation between Bogoliubov transformations and quadratic Hamiltonians and finally prove Theorems 2.15 and 2.18. Indeed, as explained in Section 2.6, Bach, Lieb and Solovej [15] introduced a general definition of quadratic Hamiltonians on fermionic Fock spaces. This refers to Definition 2.10. It leads to a strongly continuous one-parameter group of Bogoliubov transformations, as explained in the proof of Lemma 2.12. Here, we study such kind of groups and discuss their generators in relation with the Shale-Stinespring condition (see Theorem 2.8).
We take $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ and recall that a Bogoliubov transformation is a unitary operator $\mathcal{U}\in\mathcal{B}(\mathcal{H})$ such that $\mathcal{U}\mathfrak{A}=\mathfrak{A}\mathcal{U}$, as explained in Section 2.4. We thus start by relating these properties with the generator of a group of Bogoliubov transformations:
@@ 4 Technical Results: BLS’s Approach to Quadratic Operators
Proof. Any bounded operator $\mathcal{U}$ acting on $\mathcal{H}$ can always be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ by
$\mathcal{U}=\left(\begin{array}[]{cc}P_{\mathfrak{h}}\mathcal{U}P_{\mathfrak{h}}&P_{\mathfrak{h}}\mathcal{U}P_{\mathfrak{h}}^{\perp}\\ P_{\mathfrak{h}}^{\perp}\mathcal{U}P_{\mathfrak{h}}&P_{\mathfrak{h}}^{\perp}\mathcal{U}P_{\mathfrak{h}}^{\perp}\end{array}\right)$ (104)
with $P_{\mathfrak{h}}$ being the orthogonal projection on $\mathfrak{h}\oplus\{0\}$ and $P_{\mathfrak{h}}^{\perp}\doteq\mathbf{1}-P_{\mathfrak{h}}$. (By a slight abuse of notation, we can use in an appropriate way the identification $\mathfrak{h}\oplus\{0\}\equiv\mathfrak{h}$ and $\{0\}\oplus\mathfrak{h}\equiv\mathfrak{h}$ to see $P_{\mathfrak{h}}\mathcal{U}P_{\mathfrak{h}}$, $P_{\mathfrak{h}}\mathcal{U}P_{\mathfrak{h}}^{\perp}$, $P_{\mathfrak{h}}^{\perp}\mathcal{U}P_{\mathfrak{h}}$ and $P_{\mathfrak{h}}^{\perp}\mathcal{U}P_{\mathfrak{h}}^{\perp}$ as four operators acting on $\mathfrak{h}$.) If $\mathcal{U}$ is a Bogoliubov transformation, i.e., $\mathcal{U}\mathcal{U}^{\ast}=\mathcal{U}^{\ast}\mathcal{U}=\mathbf{1}$ and $\mathcal{U}\mathfrak{A}=\mathfrak{A}\mathcal{U}$ then, using (21) together with $\overline{X}\doteq\mathcal{C}X\mathcal{C}$ (Remark 2.1) one computes that the above representation can be reduced to
$\mathcal{U}=\left(\begin{array}[]{cc}\mathbf{u}&\mathbf{v}\\ \overline{\mathbf{v}}&\overline{\mathbf{u}}\end{array}\right)\ ,\qquad\mathbf{u},\mathbf{v}\in\mathcal{B}\left(\mathfrak{h}\right)\ .$ (105)
Applied to the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations, we get Assertion (i). It remains to represent the self-adjoint Hamiltonian $h$ leading to the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$. A priori, one can represent $h$ on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ like (104) for $\mathcal{U}=h$, but as explained after Equation (38), there are nontrivial domain issues. To have some control on the domain of the corresponding operators, we first use Lemma 4.1 to deduce that $h$ is a self-dual Hamiltonian, i.e., $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$. Secondly, recall that basis projections are orthogonal projections $P\in\mathcal{B}(\mathcal{H})$ satisfying $\mathfrak{A}P\mathfrak{A}=P^{\bot}\doteq\mathbf{1}_{\mathcal{H}}-P$. If $\dim\ker h\in 2\mathbb{N}\cup\{\infty\}$ and $h=-\mathfrak{A}h\mathfrak{A}$, then [18, Lemma 4.8 and its proof] leads to the existence of a pair of basis projections $P$ and $P^{\bot}\doteq\mathbf{1}-P$ with range $\mathfrak{h}_{P}$ and $\mathfrak{h}_{P^{\bot}}$, respectively, such that $h$ can be represented on $\mathcal{H}=\mathfrak{h}_{P}\oplus\mathfrak{h}_{P^{\bot}}$ by
@@
where $\mathbb{A}=\mathbb{A}^{\ast}$ is a (possibility unbounded) self-adjoint operator with domain $\mathcal{D}\left(\mathbb{A}\right)\subseteq\mathfrak{h}_{P}$. Even in the finite dimension situation, one can use the identification $\mathfrak{h}_{P}\equiv\mathfrak{A}\mathfrak{h}_{P}=\mathfrak{h}_{P^{\bot}}$ and see $\mathbb{A}^{\top}$ as an operator acting on $\mathfrak{h}_{P^{\bot}}$. We meanwhile observe from [18, Lemma 4.8] that
@@
In the above matrix representation of $h$, note that we used $X^{\top}\doteq\mathcal{C}X^{\ast}\mathcal{C}$ (Remark 2.1), which yields $\mathbb{A}^{\top}=\overline{\mathbb{A}}=\mathcal{C\mathbb{A}C}$. By [18, Lemma 3.6], there is a Bogoliubov transformation $\mathcal{U}$ such that $h$ can be represented on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$ via the following product
$h=\mathcal{U}^{\ast}\left(\begin{array}[]{cc}\mathbb{A}&0\\ 0&-\mathbb{A}^{\top}\end{array}\right)\mathcal{U}\ .$
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 4.1 and 4.2 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. Let $c\in\mathbb{R}^{+}\cup\{\infty\}$ and assume that
@@
Observe that Assumption 1 and Condition (123) imply that $\mathcal{Y}$ must be a core for the self-adjoint operator $\Upsilon_{0}$, by [44, Proposition 6.3]. In other words, $\Upsilon_{0}$ restricted to the subspace $\mathcal{Y}\subseteq\mathcal{D}\left(\Upsilon_{0}\right)$ is essentially self-adjoint. Note also that Lemma 4.3 can always be applied whenever the off-diagonal operator-valued coefficient $\overline{D}_{0}$ is relatively bounded with respect to the diagonal one $\Upsilon_{0}$ (see (44)). This is a direct consequence of the following lemma:
@@
From now, we restrict our focus from (possibly) fully unbounded cases to bounded off-diagonal operator-valued coefficients, i.e., $D_{0}\in\mathcal{B}(\mathfrak{h})$. It is worth mentioning that the diagonal element $\Upsilon_{0}$ is still taken as a possibly unbounded operator. In this case, using very standard arguments with Lemmata 4.3 and 4.4, we easily write all operator-valued coefficients of groups of Bogoliubov transformations via Dyson series involving operator-valued coefficients of their generator.
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and use Lemmata 4.1 and 4.2 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations as